The weights of steers in a herd are distributed normally. The standard deviation is 200lbs and the mean steer weight is 1000lbs. Find the probability that the weight of a randomly selected steer is between 720 and 1160 lbs. Round your answer to four decimal places.
Answer #1

Here, μ = 1000, σ = 200, x1 = 720 and x2 = 1160. We need to compute P(720<= X <= 1160). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z1 = (720 - 1000)/200 = -1.4
z2 = (1160 - 1000)/200 = 0.8

Therefore, we get
P(720 <= X <= 1160) = P((1160 - 1000)/200) <= z <= (1160 - 1000)/200)
= P(-1.4 <= z <= 0.8) = P(z <= 0.8) - P(z <= -1.4)
= 0.7881 - 0.0808
= 0.7073